A sample of trichlorine tetraoxide has a mass of 73.28 g. In theory how many moles of chlorine gas (Cl2) could be extracted from this?
1
Expert's answer
2017-02-24T14:41:05-0500
2Cl₃O₄ -> 3Cl₂

2Mol 3Mol

V(Cl₃O₄)=73.28/170.5=0.43Mol

V(Cl₂)=0.43*3/2=0.64Mol